OpenStudy (anonymous):

please help me solve [n+2]=4 [x+3]=-2 [3d+5]<-2 4[k+5]>8

OpenStudy (anonymous):

Are you using the "[" and "]" as brackets, like parentheses or are you looking for "greatest integer" or perhaps some other meaning with using brackets?

OpenStudy (anonymous):

i know im not supposed to use the brackets. usualy instead of the brackets its just a straight line but i cant find it on the keyboard

OpenStudy (anonymous):

Do you mean |n + 2| = 4 for your first question? If, you do, it's above the backslash on the far right side.

OpenStudy (anonymous):

hmmmmmmm it seems you are right lol and yes that is what i ment for all of them

OpenStudy (anonymous):

|n + 2| = 4 means BOTH n + 2 = 4 AND n+ 2 = -4, so you will get 2 values for "n". So, just solve these last 2 equations. That will give you your answer for the first one. For the second one, you use the same relationship, that is, |m| = n means m = n and m = -n. Show me you are getting this by re-writing the second one like I showed you for the first.

OpenStudy (anonymous):

x+3=2 and x+3=-2

OpenStudy (anonymous):

Very Good! So, just as you did for the first one, solve these 2 equations and you will get 2 numbers, both together are the answer. It's a 2-number answer.

OpenStudy (anonymous):

For the third one, |3d + 5| < -2 This one cannot be solved as written because the || sign makes the expression positive and you can't have a positive number less than a negative number.

OpenStudy (anonymous):

oh ok thank you so much you are an amazing tutor

OpenStudy (anonymous):

what about 4 i dont understand that one

OpenStudy (anonymous):

Still there?

OpenStudy (anonymous):

4|k + 5| > 8 means |k + 5| > 2 which means k + 5 > 2 and k + 5 < -2

OpenStudy (anonymous):

oh ok thank you that makes more sence